OpenStudy (anonymous):

Please check my answers and let me know if i got any wrong. Thank you! Convert each log statement to exponential form and each exponential statement to log form. (a) 5^-3= 1/125 Log(base5)(1/125)=-3 (b) ln2=.69315 2=e^.69315 (c ) 10^½ =3.16228 ½log3.16228 (d) e^A+B=C Log(base e)C=A+B (e) 3^0=1 log(base 3)1=0

jimthompson5910 (jim_thompson5910):

(a) and (b) are correct

jimthompson5910 (jim_thompson5910):

(c) is not correct (but there may be a typo)

jimthompson5910 (jim_thompson5910):

did you mean to say 1/2 = log 3.16228 ??? for (c)

OpenStudy (anonymous):

no typo on c….that was the one i was stuck on

OpenStudy (anonymous):

would 1/2 = log 3.16228 be the right answer for c?

jimthompson5910 (jim_thompson5910):

yes that's correct

OpenStudy (anonymous):

did you happen to look at d & e?

jimthompson5910 (jim_thompson5910):

for d, is A+B the exponent?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

then you are correct you can also say ln(C) = A+B since ln(x) = log(base e)(x)

jimthompson5910 (jim_thompson5910):

and part e) is correct

OpenStudy (anonymous):

thank you for your help!!

jimthompson5910 (jim_thompson5910):

you're welcome